A cable hangs in a parabolic arc between two poles 100 feet apart. The poles are 30 feet high and the lowest point on a suspended cable is 5 feet above the ground

a) find the equation of the arc if the vertex is the lowest points of the cable

b) find the height of the cable at a point 10 feet from one of the poles



1
Expert's answer
2021-11-10T18:25:49-0500

a) Let point "(0, 5)" be the vertex of the parabola. Then the equation of the parabola will be


"y=ax^2+5, a>0"

Given "y(-50)=y(50)=30." Substitute


"a(50)^2+5=30"

"a=0.01"

The equation of the arc is

"y=0.01x^2+5"


b)

If "x_1=50-10=40"

"y=0.01(40)^2+5"

"y=21"

The height of the cable at a point 10 feet from one of the poles is "21" feet.